Tickets for a piano concert sold for $4.50 for each adult. Student tickets sold for $2.00 each. The total receipts for 1720 tickets were $5980. Find the number of adult tickets sold.
Answer by rfer(10432)

a+s=1720, a=1720-s
4.50a+2.00s=5980
------------------
4.50(1720-s)+2.00s=5980
7740-4.5s+2s=5980
2.5s=1760
s=704
------------------
s=704*$2.00=1408
a=1016*$4.50=4572
total=$5980